Write an equation and solve.
The width of Henry's pool is 8 yards less than its length. If the area of the pool is 200 square yards, what are its length and width?

Answer by mananth(16946)

let the length of the pool be x yards
the width will be x-8
..
The area of rectangje = L*W
x(x-8)=200
x^2-8x=200
x^2-8x-200=0
..
x1, x2 the roots of the equation will be:
x1 = (8+sqrt(64+800)) / 2a
x1=18.7 yards
x2 =(8-sqrt(64+800)) / 2a
= -10.7 it is negative.
..
SO the dimensions are 18.7 & 10.7 yards